### v1

# Bimanual Regrasp Planning and Control for Eliminating Object Pose Uncertainty

Abstract
Precisely grasping an object is a challenging task due to pose uncertainties. Conventional methods have used cameras and fixtures to reduce object uncertainty. They are effective but require intensive preparation, such as designing jigs based on the object geometry and calibrating cameras with high-precision tools fabricated using lasers. In this study, we propose a method to reduce the uncertainty of the position and orientation of a grasped object without using a fixture or a camera. Our method is based on the concepts that the flat finger pads of a parallel gripper can reduce uncertainty along its opening/closing direction through flat surface contact. Three orthogonal grasps by parallel grippers with flat finger pads collectively constrain an object’s position and orientation to a unique state. Guided by the concepts, we develop a regrasp planning and admittance control approach that sequentially finds and leverages three orthogonal grasps of two robotic arms to eliminate uncertainties in the object pose. We evaluated the proposed method on different initial object uncertainties and verified that the method have satisfactory repeatability accuracy. It outperforms an AR marker detection method implemented using cameras and laser jet printers under standard laboratory conditions.
Index Terms:
Grasping Uncertainty, Bimanual Manipulation
I Introduction
Asignificant challenge in robotic manipulation lies in addressing the uncertainties associated with object grasping. The uncertainties often arise from errors in environmental registration, inaccuracies in object pose recognition, and unbalanced contact during grasping that leads to pose deviations. The uncertainties can result in discrepancies between the actual and expected pose of objects or tools, potentially causing task failures. Existing methods used to address uncertainty problems typically fall into two categories: Sensor-based correction methods [1][2] and fixture-based constraint methods [3][4]. Sensor-based methods leverage additional hardware, such as force sensors or tactile sensors, to detect and mitigate uncertainties. Fixture-based methods, on the other hand, rely on structuring the environment or using special-purpose structures to constrain object pose. However, both approaches have notable limitations, as they often require additional hardware adjustments or modifications to the environment.
Figure 1: (a) A workpiece is difficult to be accurately grasped without fixture-based positioning. (b) Flat finger pads of parallel grippers can help reduce uncertainty in the hand opening/closing direction. After grasping, the object’s uncertainty is constrained within the gray plane shown in the subfigure. (c) Three orthogonal grasps by parallel grippers constrain an exact pose. This paper proposes a bimanual regrasp planning method that generates regrasping motion sequence by consecutively executing three orthogonal grasps to eliminate object pose and grasp uncertainty.
This paper proposes a novel bimanual regrasp planning method that leverages flat finger pad constraints and admittance control to eliminate uncertainties. The proposed approach does not require external fixtures or additional sensors, making it more flexible for diverse robotic manipulation scenarios. The fundamental concepts of the proposed method are illustrated in Fig. 1(b) and (c). It is based on the observation that the flat finger pads of parallel grippers can reduce uncertainty along the hand’s opening/closing direction through flat surface contact. Three such grasps collectively constrain the object’s position and orientation to a unique state. Guided by the concepts, we develop a regrasp planning approach [5] that sequentially executes a series of regrasping motions using two robotic arms while selecting mutually orthogonal grasp triples to simultaneously eliminate uncertainties in the object pose. However, grasp planning alone does not fully achieve our goal, as simply executing the planned motion sequence using position control may exert excessive grasping forces on the uncertain object and lead to deformation or even damage. To address this issue, we develop an admittance control [6] method that utilizes force/torque sensors embedded in the robot arm’s wrists to adjust the grasping pose and conform to the uncertain object pose. By integrating regrasp planning with admittance control, our method iteratively eliminates uncertainties along three orthogonal axes, thus achieving an accurate object pose estimation.
In the experiments, we implemented the proposed method and evaluated its performance on different initial object uncertainties. The results demonstrate that the method achieves satisfactory accuracy without the need for external fixtures or additional sensors. Furthermore, we compared the accuracy of the proposed method with that of AR-based detection. The results indicate that our method outperforms the AR marker detection implemented using cameras and laser jet printers under standard laboratory conditions. The proposed approach leverages the robot’s inherent geometric constraints and force sensors to eliminate uncertainties in the grasped object. It improves task success rates by addressing uncertainties in a systematic and hardware-efficient manner.
II Related Work
The primarily related work of this paper focuses on regrasp planning and the elimination of grasping errors. Regrasp planning aims at automatically generating grasp sequences to accomplish object manipulation. It involves two main categories: One aims to identify discrete pick-and-place sequences. For example, Liu et al. [7] developed regrasp planning using multiple mobile robots. The method achieves flipping manipulation with the help of regrasping. Kim et al. [8] developed a randomized planning method for regrasp planning. Baek et al. [9] considered non-prehensile grasping in regrasp planning and secured power grasp for stable holding. Qin et al. [10] presented a regrasp planning method for installing a deformable linear object. Rigidity constraints were considered when planning regrasp motion. Mitrano et al. [11] studied a similar problem and proposed considering avoiding environment-grasp-body loops to accelerate grasp selection and planning. Sundaralingam et al. [12] extended regrasp planning to multi-fingered hand. Murooka et al. [13] proposed the foot-step and regrasping planning method for a humanoid robot to simultaneously determine foot steps and hand switch. The other focuses on continuous adjustments. For instance, Maximo et al. [14] presented the grasp space method for continuous regrasp (rolling or sliding) that maintains force closures. Hu et al. [15] formulated regrasp planning as a multi-modal problem, where three manifolds, including the regrasp manifold, the transfer manifold, and the slide manifold, were planned across to determine action sequences and thus plan motion. Cheng et al. [16] further proposed the contact mode method to guide manipulation planning across multiple manifolds. The work presented in this paper is based on iterative pick-and-place actions using two robot hands, and thus it should be categorized under the first type of regrasp planning.
One of the most critical challenges in regrasp planning is the planning speed. Researchers have explored various methods to enhance efficiency. For example, Cho et al. [17] developed a fast regrasp planning algorithm using look-up tables and two-finger parallel grippers. Chiu et al. [18] proposed a reinforcement learning method to accelerate the planning of regrasp poses. Their actions were defined in the local frame of a robot hand, thereby resolved the dependencies on robotic kinematics. Wada et al. [19] significantly improved planning speed by using a neural network to filter waypoint poses and quickly select grasp poses. Xu et al. [20] proposed learning reorientation object poses and then plan regrasp motion while considering the learned middle reorientation poses. Learning-based methods have demonstrated effectiveness in improving regrasp planning efficiency. While there are other numerous publications, this paper does not focus heavily on this aspect and thus will not delve into further details.
Another critical challenge in regrasp planning is uncertainty. Repeated regrasping can result in error accumulation, causing the grasp pose to deviate significantly from the planned outcome. The deviation can prevent the robot from successfully executing subsequent tasks. To address the uncertainty problems, Bauza et al. [21] integrated tactile information into regrasp planning and achieved precise and robust pick and placement for regrasping. Bronars et al. [22] proposed a simultaneous tactile estimation and control method to continuously plan sliding regrasp motion while maintaining grasping precisions. Gualtieri et al. [23] estimated the quality of grasp candidates against uncertainty based on point cloud completion, and optimized regrasp sequences to be robust. Zhang et al. [24] used reinforcement learning to learn a regrasp policy that optimizes grasp stability. Both vision and visual, tactile information were used for learning. This study can be partially classified as research aimed at reducing uncertainties in grasping poses during the regrasp planning process. However, our primary goal is not to achieve precise posse estimation for regrasping. Instead, we focus on actively leveraging the consecutive geometric constraints inherent in the regrasping process [25][26][27][28] to eliminate errors.
III Regrasp Planning Using Orthogonal Triples
The essence of using regrasp to eliminate uncertainty lies in leveraging the parallel finger pads of a gripper to constrain an object’s pose to a plane. By performing three successive regrasp using parallel-finger grippers along mutually orthogonal directions, an object can be constrained to three mutually orthogonal planes, and thus be precisely located. Therefore, the first step to achieve the goal of this study is to identify grasp poses along three mutually orthogonal directions. We are going to introduce the grasp planning and grouping methods in the first subsection below. Once the grasp poses and groups are determined, search and motion planning techniques can be employed to generate regrasp actions and consequently the robot motion required to eliminate uncertainties. The second subsection below details the specific methods for search and motion planning.
III-A Orthogonal Grasp Triples
[…]
Figure 2: (a) Antipodal grasp planning. (b) Candidate grasp poses planned using the antipodal method. (c) Separating the candidate grasp poses into three mutually orthogonal groups considering their opening and closing directions. Each color in the subfigure represents one group.
Subsequently, based on the results of the grasp planning, we proceed by separating them into groups. Since our goal is to identify three mutually orthogonal grasp poses, we set our grouping criterion to be based on the opening and closing directions of the gripper fingers. Grasp poses with the same opening and closing direction are grouped together, and each group can be represented by a unique opening and closing vector, denoted as $\boldsymbol{v}_{i}$ Note 1 Note 1 Note 1In this context, the direction of $\boldsymbol{v}_{i}$ is not considered. That is, $\boldsymbol{v}_{i}$ assumed identical to $-\boldsymbol{v}_{i}$.. Through the grouping process, we can obtain several groups represented by $\boldsymbol{v}_{1}$, $\boldsymbol{v}_{2}$, …The candidate grasp poses in Fig. 2(b) can be right classified into three mutually orthogonal groups. They are visualized in Fig. 2(c) using red, green, and blue colors.
After obtaining the grasp groups, we perform triple combinations of the groups to identify three groups that satisfy the mutually orthogonal constraint. To evaluate whether the constraint is satisfied, we use the following formula to compute a score for each triple combination.
$\mathtt{score}=|\boldsymbol{v}_{i}\cdot\boldsymbol{v}_{j}|+|\boldsymbol{v}_{i}\cdot\boldsymbol{v}_{k}|+|\boldsymbol{v}_{j}\cdot\boldsymbol{v}_{k}|$ (1)
The $\boldsymbol{v}_{i}$, $\boldsymbol{v}_{j}$, and $\boldsymbol{v}_{k}$ in the above equation represent the opening and closing directions of three different groups in a triple. We sort the scores of all triples and prioritize them with a score closest to 0 for regrasp planning and uncertainty elimination. The closer the score value is to 00, the more orthogonal the three grasps from the groups are to each other. By performing successive regrasp using the prioritized triples, object uncertainties can be effectively minimized. The example shown in Fig. 2 represents a special case where the planned grasp poses are right divided into three orthogonal groups with a score of zero. In more general cases, there may be multiple candidate groups and triple combinations. The triples may have non-zero scores and regrasp should be performed by prioritizing the ones with smaller scores.
III-B Regrasp Planning
For the sorted triples, motion planning is sequentially performed on each of them, and the first successful solution is output as a valid candidate. Note that each triple comprises three groups of grasp poses. During planning, one pose from each group needs to be selected to form a combination of three grasps for computation. In this approach, we do not explicitly extract the three-grasp combinations as a separate step. Instead, we leverage the two arms of a bimanual robot to perform regrasp planning interactively for the three grasp groups. The grasps in each group are sequentially assigned to the two robotic arms for planning. In this way, we incrementally explore the combinations of the grasp triples, and output successful motion once found. The detailed algorithmic flow is illustrated in Fig. 3. Note that the process illustrated in the figure is merely an incremental search. If a large amount of data can be accumulated using the proposed method in the future, we strongly recommend exploring learning-based methods to improve efficiency.
Figure 3: Regrasp planning used in this study is a simple incremental search method, which sequentially traverses grouped grasp triples until a valid solution is found. Since the regrasp planning algorithm itself is not the primary focus of this paper, further optimization aspects such as acceleration or grasp selection refinement are not explored.
IV Conforming to Object Pose Uncertainty
After planning a feasible sequence of bimanual robot motions, the next step is execution. However, due to uncertainties in the object’s pose, a significant discrepancy may exist between the real-world pose and the planning simulation, making pure position control infeasible. This subsection discusses this limitation and highlights the necessity of conforming to object pose uncertainty. It then introduces the admittance control method and explains how it enables the system to adapt to uncertain object poses during handover regrasping.
Consider the scenario shown in Fig. 4(a), where the left hand has already moved the object from its initial pose to the handover pose, and the right hand is approaching to perform the regrasp. The target pose of the right hand is planned based on an assumed pose in the simulation. However, due to uncertainties in the real world, the object’s assumed pose in the simulation deviates from its actual pose, as illustrated by the orange dashed frames and red solid frames in Fig. 4(a). Consequently, the right hand’s grasping pose becomes misaligned with the actual object pose. Applying position control for grasping in this scenario may lead to undesirable outcomes. In mild cases, as shown in Fig. 4(b) and (c), the object may undergo deformation. In more severe cases, it may sustain damage. Therefore, rigidly executing the planned grasping motion solely with position control is not advisable.
Figure 4: (a) The right hand approaches to grasp the object. Due to uncertainties, the object is not positioned exactly at the center of the two fingers as assumed but is instead slightly shifted upward. (b) At this point, if position control is used, the upper-side finger will make contact with the object first, exerting excessive force on it and causing deformation. (c) Comparison of object deformation before and after grasping.
To address this issue, this study employs admittance control during robotic grasping and enables the manipulator’s pose to conform to the object’s pose. The basic concept of admittance control is to establish a mass-damper-spring system’s motion equation based on the external force exerted on the robot and the desired target force . The robot then follows the solution of this equation to adjust its motion accordingly, thereby accommodating external forces. For a desired robot position $x_{d}$, the admittance control equation is
$M(\ddot{\mathit{x}}_{p}-\ddot{\mathit{x}}_{d})+B(\dot{\mathit{x}}_{p}-\dot{\mathit{x}}_{d})+K(\mathit{x}_{p}-\mathit{x}_{d})=\mathit{f}+\mathit{f}_{d}.$ (2)
Here, $x_{p}$, $\dot{x_{p}}$, and $\ddot{x_{p}}$ are respectively the measured current robot end-effector position, velocity, and acceleration. $M$, $B$, $K$ denote the virtual inertia, damping, and stiffness parameters. $f$ is the force applied to the system, while $f_{d}$ represents the target force exerted by the robot after applying admittance control.
In our specific grasping scenario, the target force is set to zero, as our objective is to ensure that the robot does not exert any external force on the object beyond the necessary grasping force, thereby preventing unintended movement, deformation, or damage. During the grasping process, when one of the robotic fingers first makes contact with the object, it receives a reaction force from the object, which corresponds to the external force $f$ in equation (2). Based on the robot’s current $x_{p}$, $\dot{x_{p}}$, $\ddot{x_{p}}$, as well as the force $f$ measured using a force/torque sensor installed at the robot manipulator’s wrist, we can compute the desired position $x_{d}$, velocity $\dot{x_{d}}$, and acceleration $\ddot{x_{d}}$ using the equation. These computed values are then used to control the robot, enabling it to conform to the object’s uncertain pose during the grasping process.
Fig. 5 illustrates the effect of applying admittance control to the grasping process shown in Fig. 4. In Fig. 5(a), the upper-side finger first makes contact with the object and experiences a reaction force $f$. The admittance control then adjusts the robotic hand by moving it in a direction that balances this force, as shown in Fig. 5(b). The adjustment continues until both fingers are in contact with the object and their grasping forces reach equilibrium. At this point, $f$ disappears and the grasping forces become internal forces within the whole gripper-object system. The final grasping result is depicted in Fig. 4(c). Unlike position control, admittance control ensures that the object’s pose remains unchanged after grasping.
Figure 5: Admittance control helps avoid deformation or damage. (a) When the upper-side finger makes contact with the object, a reaction force $f$ is detected. (b) Admittance control adjusts the robot hand by moving it in a direction that balances $f$. (c) After admittance control, the object is firmly grasped, and no deformation occurs in its shape.
V Pose Estimation Using Conformed Grasp Poses
Next, we present the method for estimating the object’s pose. In the planning simulation, we used an assumed object pose. This pose serves as the basis for all grasping and planning processes. Since there is no uncertainty in the simulation, the final object pose remains deterministic. Here, we denote the object’s final pose in the planning simulation as $\boldsymbol{p}_{o_{s}}$ and $\boldsymbol{R}_{o_{s}}$, and denote the three consecutive grasping poses of the robot in the simulation as $\boldsymbol{g}_{s_{1}}$, $\boldsymbol{g}_{s_{2}}$, and $\boldsymbol{g}_{s_{3}}$. Each $\boldsymbol{g}_{s_{i}}$ consists of ($\boldsymbol{p}_{s_{i}}$, $\boldsymbol{R}_{s_{i}}$, $d_{s_{i}}$), where $\boldsymbol{p}_{s_{i}}$ and $\boldsymbol{R}_{s_{i}}$ represent the position and rotation of the grasping center point, and $d_{s_{i}}$ denotes the jaw opening width. On the other hand, by executing the planned motion in the real world, we obtain three consecutive grasping poses of the bimanual robot in the real world as $\boldsymbol{g}_{r_{1}}$, $\boldsymbol{g}_{r_{2}}$, and $\boldsymbol{g}_{r_{3}}$. Each $\boldsymbol{g}_{r_{i}}$ similarly consists of ($\boldsymbol{p}_{r_{i}}$, $\boldsymbol{R}_{r_{i}}$, $d_{r_{i}}$). The first grasping pose $\boldsymbol{g}_{r_{1}}$ aligns exactly with the planned result, while the second grasping pose $\boldsymbol{g}_{r_{2}}$ and the third grasping pose, $\boldsymbol{g}_{r_{3}}$, are adjusted through admittance control to conform to the object’s actual pose. They are different from $\boldsymbol{g}_{s_{2}}$ and $\boldsymbol{g}_{s_{3}}$. Our objective is to estimate the real-world object pose $\boldsymbol{p}_{o_{r}}$ and $\boldsymbol{R}_{o_{r}}$ based on these known values.
To address this issue, let us first consider the simulation scenario. Since our three grasping poses involve planar contacts and are mutually orthogonal, the object constrained by them should have a unique pose. The pose can be computed as follows. For each grasp, we can obtain a corresponding opening and closing vector $\boldsymbol{v}_{i}=\boldsymbol{R}_{s_{i}}(\boldsymbol{y})$. Here, $\boldsymbol{R}_{s_{i}}(\boldsymbol{y})$ denotes the $\boldsymbol{y}$ axis (2nd column) of the rotation matrix. We assume the robot hand’s opening and closing direction is aligned with its local $\boldsymbol{y}$ axis. Since the three grasping poses are mutually orthogonal, they uniquely determine the following rotation matrix
$\boldsymbol{R}_{g_{s}}=\begin{bmatrix}\boldsymbol{R}_{s_{1}}(\boldsymbol{y})&\boldsymbol{R}_{s_{2}}(\boldsymbol{y})&\boldsymbol{R}_{s_{3}}(\boldsymbol{y})\end{bmatrix}.$ (3)
Meanwhile, for each grasp, its two parallel fingers define a midplane. Given that the three grasps are mutually orthogonal, their corresponding midplanes intersect at a unique intersection point. This intersection represents a unique position constraint imposed by the grasping configuration. It can be computed using
$\boldsymbol{p}_{g_{s}}=(\boldsymbol{R}_{g_{s}}^{\top})^{-1}\begin{bmatrix}\boldsymbol{R}_{s_{1}}(\boldsymbol{y})\cdot\boldsymbol{p}_{{s_{1}}}\\ \boldsymbol{R}_{s_{2}}(\boldsymbol{y})\cdot\boldsymbol{p}_{{s_{2}}}\\ \boldsymbol{R}_{s_{3}}(\boldsymbol{y})\cdot\boldsymbol{p}_{{s_{3}}}\end{bmatrix}$ (4)
The ($\boldsymbol{p}_{g_{s}}$, $\boldsymbol{R}_{g_{s}}$) obtained from the above two equations are the unique pose determined by the three grasping poses in the planning simulation. Note that it is not the object’s final pose ($\boldsymbol{p}_{o_{s}}$, $\boldsymbol{R}_{o_{s}}$). However, since both are uniquely determined. There is a unique offset between them written as
$\boldsymbol{p}_{\text{diff}}=\boldsymbol{p}_{o_{s}}-\boldsymbol{p}_{g_{s}},~{}\boldsymbol{R}_{\text{diff}}=\boldsymbol{R}_{g_{s}}^{\top}\boldsymbol{R}_{o_{s}}$ (5)
Fig. 6 illustrates the three planes formed by three grasping poses and their intersection point.
Figure 6: Three orthogonal grasping poses with planar contacts determine three planes that intersect at a unique point. Thus, a unique position $\boldsymbol{p}_{g_{s}}$ is determined. Meanwhile, the opening-closing directions of the three gasping poses help constrain a unique rotation $\boldsymbol{R}_{g_{s}}$
In the real world, $\boldsymbol{p}_{o_{r}}$ and $\boldsymbol{R}_{o_{r}}$ are uniquely constrained by the grasping poses and have the same offset from the pose determined by the grasping poses. That is
$\boldsymbol{p}_{o_{r}}=\boldsymbol{p}_{g_{r}}+\boldsymbol{p}_{\text{diff}},~{}\boldsymbol{R}_{o_{r}}=\boldsymbol{R}_{g_{r}}\boldsymbol{R}_{\text{diff}}.$ (6)
Thus, $\boldsymbol{p}_{o_{r}}$ and $\boldsymbol{R}_{o_{r}}$ can be computed. Here, $\boldsymbol{p}_{g_{r}}$ and $\boldsymbol{R}_{g_{r}}$ can be obtained similarly to equations (4) and (3) using ($\boldsymbol{p}_{r_{1}}$, $\boldsymbol{R}_{r_{1}}$), ($\boldsymbol{p}_{r_{2}}$, $\boldsymbol{R}_{r_{2}}$), and ($\boldsymbol{p}_{r_{3}}$, $\boldsymbol{R}_{r_{3}}$).
VI Experiments and Analysis
This section presents the real-world validation of the proposed method. Our experimental platform consists of two UR3e robots from Universal Robots. Their installation setup is shown in Fig. 7. The UR3e series robots are equipped with high-precision force/torque sensors at their end effectors, which facilitate the implementation of admittance control. Each robot is installed with a Robotiq Hand-E gripper from Robotiq. The opening range of the gripper is 0$\sim$50 mm. The regrasp planning is performed using the WRS simulation environment Note 2 Note 2 Note 2https://github.com/wanweiwei07/wrs. Fig. 6 in the previous section provided a view of the simulation environment. The coordinate system of our experimental platform is defined with its origin at the intersection of the central plane between the two robot arms and the rear edge of the table, as illustrated by the colored frame in Fig. 7. The $\boldsymbol{x}$-axis points forward, the $\boldsymbol{y}$-axis points towards the left arm, and the $\boldsymbol{z}$-axis aligns with the normal direction of the tabletop surface.
[…]
To ensure clarity in comparison and analysis, we used the simple L-shaped object shown in Fig. 7 and assumed a single initial pose in the planning simulation. The object has a long edge of $125$ mm and a short edge of $100$ mm, with a square cross-section of $25$ mm $\times$ $25$ mm. The surfaces of the L-shaped object are mutually perpendicular. They could align with the three axes of the coordinate system and facilitate analysis. The initial pose of the object in the simulation was set to $(700,300,1.25)$ mm. Its rotation was set to be aligned with the coordinate frame of the robot platform (recall the colored frame shown in Fig. 7). The long edge aligns with the $\boldsymbol{y}$ axis. The short edge aligns with the $\boldsymbol{z}$ axis. Similarly to the initial pose, we set the regrasp pose of the object in the simulation to be at a unique fixed location $(680,50.0,120.0)$ mm, with the rotation identical to the initial one to facilitate analysis. The location is in the middle of the two arms, making it easy to plan effective bimanual regrasp motions. The regrasp motion planned based on the parameter settings and the proposed planning methods led to orthogonal triples where the first selected grasp is along the $\boldsymbol{y}$ axis, the second grasp is along the $\boldsymbol{z}$ axis, and the third grasp is along the $x$ axis. Together, the three grasps uniquely constrain the object’s pose.
In the real world, we manually introduced uncertainties in multiple directions to the object for experimental evaluation. These uncertainties are illustrated in Fig. 8. Fig. 8(a) represents initial object poses obtained by introducing uncertainties in the horizontal directions (back, front, left, and right) relative to a center pose shown in Fig. 8(a.v). The center pose closely corresponds to the initial pose of the object used in the simulation. Note that the center pose is not exactly the initial pose in the simulation. The front, back, left, and right movements are approximate ±$\pm 12.5$ mm and are also not exact. We are not enforcing precise placement of the object when preparing these poses. The reason is that preparing exact uncertainty in the real world can be challenging. Finger contact during the grasping process naturally causes additional errors. Specifying exact uncertainty values and evaluating absolute accuracy would not provide meaningful insights. We analyze repeatability errors rather than absolute errors to evaluate the effectiveness of our method in eliminating uncertainties. In addition to the horizontal uncertainties, we also prepared initial poses with uncertainties in the upward, upper-front, and upper-back directions. These poses are shown in Fig. 8(b). The L-shape object is elevated using a white block while being shifted back and forward with the same approximate ±$\pm 12.5$ mm offset when preparing these uncertain poses.
Figure 8: (a) $9$ uncertain poses in the horizontal directions. (a.i) BL – Back Left. (a.ii) BC – Back Center. (a.iii) BR – Back Right. (a.iv) L – Left. (a.v) C – Center. (a.vi) R – Right. (a.vii) FL – Front Left. (a.viii) FC – Front Center. (a.ix) FR – Front Right. (b) $3$ uncertain poses in upward directions. (b.i) DC – Down Center (for reference). (b.ii) UC – Upper Center. (b.iii) Upper Back. (b.iv) Upper Front.
Using the planned robot motions, we performed regrasping and applied admittance control following the proposed method for each initial object pose in Figs. 8(a) and (b), and estimated their final object poses using equation 6. The estimation results are summarized in the “Proposed method” columns of Table I. The results corresponding to (a) and (b) are separated by a double solid line. From the results, we can observe that the estimation performs reasonably well. In the case of the uncertain object placement shown in Figs. 8(a), the estimated $\boldsymbol{p}_{o_{r}}$ vector has a mean and variance of $(9.82,0.002)$ mm and $(898.33,0.019)$ mm in the $\boldsymbol{y}$ and $\boldsymbol{z}$ directions respectively. The low variance aligns well with our intuition – For the uncertain poses listed in Fig. 8(a), the object is placed on the table. Its position in the $\boldsymbol{z}$-direction remains the same. Meanwhile, the object is always grasped by fingers closing from the $\boldsymbol{y}$-direction. Its uncertainty in the $\boldsymbol{y}$-direction is constrained first grasp’s two fingers. Therefore, the final estimated poses should be consistent (have a low variance) in these two directions. On the other hand, the first values of the estimated $\boldsymbol{p}_{o_{r}}$ can be clearly categorized into three distinct groups. Among them, BC and BR form one group, with a mean and variance of $(687.05,2.1)$ mm. L, C, and R constitute another group, with a mean and variance of $(699.3,1.4)$ mm. Finally, FL and FC belong to the third group, with a mean and variance of $(709.45,2.1)$ mm. This result is also consistent with our initial placement. The object displacement in the $\boldsymbol{x}$ direction can be clearly distinguished, and the repeatability is satisfying. The variance of each distinct group falls within the range of the manual placement uncertainty. Similar analysis and conclusions can be drawn from the results for the uncertain object placements shown in Figs. 8(b). They are in the UB, UC, and UF rows. The estimated $\boldsymbol{p}_{o_{r}}$ vectors have nearly the same values in the $\boldsymbol{y}$ and $\boldsymbol{z}$ directions and clearly distinct values in the $\boldsymbol{x}$ direction. For rotation, all estimated results are generally aligned with the coordinate axes. They have an average deviation of $0.59^{\circ}$ with $0.11^{\circ}$ variance from rotation matrix $\begin{bmatrix}\boldsymbol{x}&\boldsymbol{z}&-\boldsymbol{y}\end{bmatrix}$. This outcome is consistent with the fact that the object’s initial placement is always on a plane and that the opening-closing direction of the first grasp is along the $\boldsymbol{y}$ axis. Its repeatability is also satisfying considering the small variance. Note that the BL and FR rows in the table do not contain values. This is because the object’s pose uncertainty was too large, causing the robot to collide with the object during execution. The maximum tolerable uncertainty in our method is constrained by the gripper’s opening width. When the uncertainty exceeds half of the gripper’s opening width, collisions may occur and lead to execution failure. Additionally, the table does not include estimation results for DC pose. This pose is only provided to give readers a reference for comparison before the object is elevated, and thus, no estimation was performed for it.
TABLE I: Pose estimation results
Type Exec. Proposed method AR marker
$\boldsymbol{p}_{o_{r}}$ (mm) $\boldsymbol{R}_{o_{r}}$ $\boldsymbol{p}_{o_{r}}$ (mm) $\boldsymbol{R}_{o_{r}}$
BL Fail. - - - -
BC Succ. $(685.6,49.8,118.2)$ $\begin{bmatrix}1.00&-0.00&-0.00\\ -0.01&0.00&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(715.4,56.5,128.3)$ $\begin{bmatrix}1.00&-0.01&0.07\\ 0.07&-0.02&-1.00\\ 0.01&1.00&-0.02\end{bmatrix}$
BR Succ. $(688.5,49.8,118.3)$ $\begin{bmatrix}1.00&-0.00&-0.00\\ -0.02&-0.00&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(719.5,34.8,146.3)$ $\begin{bmatrix}1.00&-0.06&0.00\\ 0.00&-0.02&-1.00\\ 0.06&1.00&-0.02\end{bmatrix}$
L Succ. $(699.5,49.9,118.5)$ $\begin{bmatrix}1.00&-0.01&0.00\\ 0.00&-0.00&-1.00\\ -0.00&1.00&0.00\end{bmatrix}$ $(740.2,54.1,131.5)$ $\begin{bmatrix}1.00&-0.03&0.05\\ 0.05&-0.01&-1.00\\ 0.03&1.00&-0.00\end{bmatrix}$
C Succ. $(697.8,49.9,118.1)$ $\begin{bmatrix}1.00&-0.00&0.00\\ -0.00&-0.00&-1.00\\ -0.01&1.00&-0.00\end{bmatrix}$ $(726.9,36.8,144.5)$ $\begin{bmatrix}1.00&-0.07&-0.01\\ -0.01&-0.01&-1.00\\ 0.07&1.00&-0.01\end{bmatrix}$
R Succ. $(700.7,49.8,118.3)$ $\begin{bmatrix}1.00&-0.00&-0.00\\ -0.00&-0.00&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(726.3,55.6,145.8)$ $\begin{bmatrix}1.00&-0.07&0.07\\ 0.06&-0.01&-1.00\\ 0.07&1.00&-0.01\end{bmatrix}$
FL Succ. $(710.9,49.8,118.4)$ $\begin{bmatrix}1.00&0.01&-0.00\\ 0.03&-0.00&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(741.3,57.7,128.4)$ $\begin{bmatrix}0.99&-0.01&0.10\\ 0.10&-0.01&-0.99\\ 0.01&1.00&-0.01\end{bmatrix}$
FC Succ. $(708.0,49.8,118.5)$ $\begin{bmatrix}1.00&0.00&-0.00\\ 0.02&-0.00&-1.00\\ -0.01&1.00&-0.00\end{bmatrix}$ $(756.1,29.4,110.6)$ $\begin{bmatrix}1.00&-0.00&0.01\\ 0.01&-0.03&-1.00\\ 0.00&1.00&-0.03\end{bmatrix}$
FR Fail. - - - -
UB Succ. $(689.8,49.8,131.8)$ $\begin{bmatrix}1.00&0.01&0.00\\ -0.02&-0.01&-1.00\\ -0.03&1.00&0.00\end{bmatrix}$ $(717.5,38.3,157.9)$ $\begin{bmatrix}1.00&-0.06&-0.02\\ -0.02&-0.03&-1.00\\ 0.06&1.00&-0.03\end{bmatrix}$
UC Succ. $(698.7,49.8,131.6)$ $\begin{bmatrix}1.00&-0.01&0.00\\ 0.00&-0.00&-1.00\\ -0.00&1.00&0.00\end{bmatrix}$ $(732.4,32.1,160.6)$ $\begin{bmatrix}1.00&-0.07&0.02\\ 0.02&-0.02&-1.00\\ 0.07&1.00&-0.02\end{bmatrix}$
UF Succ. $(706.6,49.8,131.6)$ $\begin{bmatrix}1.00&-0.00&-0.00\\ 0.01&-0.02&-1.00\\ -0.01&1.00&0.00\end{bmatrix}$ $(735.6,36.6,158.5)$ $\begin{bmatrix}1.00&-0.08&0.02\\ 0.01&-0.02&-1.00\\ 0.08&1.00&-0.02\end{bmatrix}$
[…]
VII Conclusions and Future Work
This paper proposes a method for eliminating object uncertainties by leveraging the grasping constraints of parallel grippers with flat finger pads and a bimanual regrasp planning method. We applied this method to estimate grasping errors for a roughly placed L-shaped object. The results demonstrate that the proposed approach effectively eliminates uncertainties and achieves high repeatability accuracy.
Currently, this method is only applicable when all three grasps involve planar contact. For curved surface contacts, additional grasps may be required to fully constrain uncertainties. Further exploring the necessary grasp combinations will be an important future topic.